# Nothing Conformal about Adaptive Conformal Inference

### v2

@@ Abstract
Conformal prediction is a widely-used framework for distribution-free uncertainty quantification, which generates valid prediction sets at a user-defined significance level. However, this framework relies on the assumption that the data-generating distribution is exchangeable, a condition that is frequently violated in time-series and other structured data. In such cases, the validity guarantees of conformal prediction break down.
Adaptive conformal inference (ACI) has been proposed as a solution for non-exchangeable data by dynamically adjusting the significance level to retain at least finite sample guarantees on the marginal coverage error rate. This paper demonstrates that, despite its name, ACI does not strictly require the use of conformal predictors. Instead, it can operate effectively with the more general concept of a confidence predictor, which is often computationally simpler. The key requirement is that larger significance levels correspond to smaller prediction sets, a property known as nested prediction sets.
Through experiments on synthetic and real-world data, we investigate whether ACI with conformal predictors offers advantages over confidence predictors. Our results indicate that confidence predictors can perform just as well, and sometimes better than conformal predictors in some cases, although further empirical studies are needed to determine when one approach may be preferable.
@@
Prediction Intervals , Coverage Guarantee , Adaptive Conformal Inference , Conformal Prediction
@@ 1 Introduction
Conformal prediction (CP) [1, 2] is a general framework for distribution-free uncertainty quantification. It can be used with essentially any machine learning algorithm and has guaranteed properties of validity, provided that the data is drawn from a probability distribution that is exchangeable, meaning essentially that any permutation is equally probable. For details on exchangeability, see e.g. [2]. The strongest notions of validity are achieved in the online setting, where Reality outputs successive examples $z_{n}:=(x_{n},y_{n})\in\boldsymbol{Z}:=\boldsymbol{X}\times\boldsymbol{Y}$, each one consisting of an object $x_{n}\in\boldsymbol{X}$ and its associated label $y_{n}\in\boldsymbol{Y}$, CP produces prediction sets, $\Gamma_{n}^{\varepsilon}$, at a user specified significance level $\varepsilon$, using $z_{1},\dots,z_{n-1}$ and $x_{n}$ as input.
While powerful and general, the validity guarantees of CP relies on the exchangeability assumption, which is a limitation e.g. in time-series forecasting. To address these limitations, many authors have suggested solutions to retain at least some validity guarantees even when exchangeability is violated. These include introducing weights to attempt to deal with distribution shifts [3, 4], and what could be described as control theoretic approaches that attempt to achieve the desired error rate by means of varying the significance level [5, 6].
Adaptive conformal inference (ACI) [5] falls into the second category, of control theoretic approaches, and has become a popular choice for non-exchangeable data. It has been suggested as a good method for time-series forecasting [7], and the software package MAPIE [8] has implemented ACI for this purpose. The more complicated conformal PID algorithm [6] can be viewed as a generalisation of ACI, as the authors recover it as a special case of conformal PID.
@@
In this paper, we will show that there is no need to use ACI with a conformal predictor. The same guaranteed error rates hold even if we use the more general notion of a confidence predictor, which can often be much more computationally tractable. Since nothing is gained by using the heavier machinery of conformal predictors in terms of error rate, the question is whether CP offers more efficient prediction sets.
@@
The rest of this paper is organised as follows. Section 2 introduces conformal and confidence predictors, as well as the relevant notation. Readers familiar with the terminology used in [2] may find it convenient to skip ahead to Section 3, where ACI and its finite sample coverage guarantee is introduced, as well as a trivial but technically necessary extension of conformal and confidence predictors. Our main result is given in Section 4, where we restate Lemma 4.1 in [5], which is the key result that is used to prove the finite sample guarantee of ACI. We indicate that this lemma does not rely on any particular property of conformal predictors, or even confidence predictors. We then argue in Section 5 that, while not strictly necessary for the finite sample guarantee of ACI, confidence predictors represent the natural way of predicting at a confidence level. Our numerical experiments on synthetic and real data are described in Section 6, and Section 7 concludes. Additional details on some algorithms used in the experiments, and a summary of the results without using ACI (i.e. at a fixed significance level) are given in the appendices.
@@ 2 Theoretical background
This section introduces the relevant theoretical background on mathematical objects and notation. Readers familiar with conformal prediction, and the notation used in [2] may find it convenient to skip ahead to the next section, where ACI is introduced.
@@ 2.1 Confidence prediction in the online setting
called examples. For notational convenience, we write $z_{i}:=(x_{i},y_{i})\in\boldsymbol{X}\times\boldsymbol{Y}:=\boldsymbol{Z}$. The measurable space $\boldsymbol{Z}$ is called the example space. Thus, the infinite sequence (1) is an element of the measurable space $\boldsymbol{Z}^{\infty}$, and we assume that it is drawn from some probability distribution $P$ on $\boldsymbol{Z}^{\infty}$. The standard assumption in CP is that $P$ is exchangeable, but in this paper we will make no such assumption, and $P$ can be any distribution.
@@
This last property is called the property of nested prediction sets.
Recall that $2^{\boldsymbol{Y}}$ is the set of all subsets of $\boldsymbol{Y}$. Thus, a confidence predictor outputs a subset of the label space based on all previous examples and the current object, and the idea is that it should contain the true label $y_{n}$ with a user specified confidence.
@@
at significance level $\varepsilon$. We say that $\Gamma$ makes an error at the $n$th trial if $y_{n}\notin\Gamma_{n}^{\varepsilon}$. More precisely,
@@ 2.2 Validity
is a sequence of independent Bernoulli random variables with parameter $\varepsilon$. In words, the event of making an error is like getting heads when tossing a biased coin, where the probability of getting heads is always $\varepsilon$. There is also the notion of conservative validity, which is more complicated to state, but essentially means that the error sequence is dominated by a sequence of independent Bernoulli variables with parameter $\varepsilon$. For the complete statement, we refer to [2].
@@ 2.3 Conformal predictors
Conformal predictors rely on the notion of a nonconformity score. Informally, this is a function that quantifies how “strange” or “unusual” an example $z$ is in relation to what we have seen before. More formally, given a sequence of examples $z_{1},\dots,z_{n}$, we form the bag (or multiset) $\Sigma_{n}:=\lbag z_{1},\dots,z_{n}\rbag$. The set of all bags of size $n$, formed from examples in $\boldsymbol{Z}$ is denoted $\boldsymbol{Z}^{(n)}$. A nonconformity measure is a measurable function
$A:\boldsymbol{Z}^{(*)}\times\boldsymbol{Z}\to\overline{\mathbb{R}}.$ (10)
@@
The main disadvantage of conformal predictors is their, often, intractable computational cost. In a regression setting, where $\ Y=\mathbb{R}$, we would theoretically have to compute the fraction in (11) for each real number, which is clearly impossible. For some special cases, such as ridge regression, efficient implementations exist, but generally the computational cost is too high. For this reason, inductive conformal predictors (ICP) were introduced. We give a brief description here, and refer the interested reader to [2] for more details. Given a training set of size $l$, we split it into two parts: the proper training set $z_{1},\dots,z_{m}$ of size $m$, and the calibration set $z_{m+1},\dots,z_{l}$ of size $l-m$. For every test object $x_{i}$, compute the prediction set
@@
These are the most widely used conformal predictors in practice, but as always, there is a price to pay for the computational efficiency. ICPs are no longer exactly valid, but their valid property are of the PAC type, with two parameters. For details on the training-conditional validity of ICPs, see [9].
@@ 3 Adaptive conformal inference
Adaptive conformal inference [5] was suggested as a way to achieve asymptotic validity under non-exchangeable data. The idea is that instead of using the same significance level for all predictions, we use the online update
@@
where $\gamma$ is a step size (or learning rate). In words, if we made an error in the last step, we increase the significance level, and if not, we decrease it. Gibbs and Candès proved the following finite sample guarantees of ACI.
@@
which in particular converges to 0 as $N\to\infty$, ensuring asymptotic validity. Note that if we want a certain absolute error deviation bound, $\delta>1/n$, for a desired error rate $\varepsilon$, with a finite sample size $N$, we can almost surely achieve it by choosing
@@
The iteration (15) could cause $\varepsilon_{t}\leq 0$ or $\varepsilon\geq 1$. Technically, prediction sets are undefined for confidence predictors, and in particular conformal predictors for $\varepsilon\notin(0,1)$. Thus, we introduce a trivial extension that is compatible with ACI.
@@
Since the difference between confidence predictors and extended confidence predictors is minor, and since any confidence predictor can be trivially modified to be an extended confidence predictor, we will use the terms interchangeably. In analogy with definition 3, we can also define an extended conformal predictor by requiring that the output is $\emptyset$ if $\varepsilon\geq 1$ and $\boldsymbol{Y}$ if $\varepsilon\leq 0$. Similar to confidence predictors and extended confidence predictors, we will use the terms conformal predictor and extended conformal predictor interchangeably.
ACI was introduced in 2021 by Gibbs and Candès [5]. Since then, it has been suggested for time-series forecasting [7], where exchangeability is typically an issue for conformal predictors. It is implemented in this mode in the python package MAPIE [10]. In the time-series domain, it has also been suggested for multi-step ahead forecasting by e.g. [11, 12], and has been shown to be a special case of the conformal PID algorithm [6]. However, despite its name, there is nothing conformal about ACI, as we shall see in the next section.
4 Nothing conformal is required
Interestingly, the proof of (16) does not use any property of a conformal, or even confidence predictor. In fact, all that is required of a predictor $\Gamma$ is that $\emptyset$ is predicted for $\varepsilon\geq 1$ and $\boldsymbol{Y}$ for $\varepsilon\leq 0$. For all other significance levels, even completely random prediction sets ensure (16). The result is proved by Lemma 4.1 in [5], which we reproduce here with notation slightly modified to align with ours.

### v4

# Beyond Conformal Predictors: Adaptive Conformal Inference with Confidence Predictors

@@ Abstract
Adaptive Conformal Inference (ACI) provides finite-sample coverage guarantees, enhancing the prediction reliability under non-exchangeability. This study demonstrates that these desirable properties of ACI do not require the use of Conformal Predictors (CP). We show that the guarantees hold for the broader class of confidence predictors, defined by the requirement of producing nested prediction sets, a property we argue is essential for meaningful confidence statements. We empirically investigate the performance of Non-Conformal Confidence Predictors (NCCP) against CP when used with ACI on non-exchangeable data. In online settings, the NCCP offers significant computational advantages while maintaining a comparable predictive efficiency. In batch settings, inductive NCCP (INCCP) can outperform inductive CP (ICP) by utilising the full training dataset without requiring a separate calibration set, leading to improved efficiency, particularly when the data are limited. Although these initial results highlight NCCP as a theoretically sound and practically effective alternative to CP for uncertainty quantification with ACI in non-exchangeable scenarios, further empirical studies are warranted across diverse datasets and predictors.
@@
Adaptive Conformal Inference (ACI) , Confidence Predictors , Non-exchangeable Data , Coverage Guarantee , Conformal Prediction (CP)
@@
{graphicalabstract}[Uncaptioned image]
{highlights}
Shows that adaptive conformal inference (ACI) achieves finite-sample coverage guarantees without needing conformal predictors (CP).
Demonstrates empirically that non-conformal confidence predictors (NCCP) can offer comparable or better performance than CP for non-exchangeable data.
Highlights that inductive NCCP avoids sacrificing training data for calibration, potentially improving predictive efficiency compared to inductive CP, especially when data are limited.
Clarifies that the primary advantage of NCCP in online settings may be computational efficiency, whereas in batch settings, it relates to improved data utilisation.
Argues that nested prediction sets are necessary for meaningful confidence-based predictions regardless of the underlying predictor type.
@@ 1 Introduction
The need for uncertainty quantification in pattern recognition arises in many safety-critical applications, such as medical diagnosis and autonomous driving, where recognising and addressing uncertainty can prevent costly or dangerous mistakes. Unfortunately, many models, including modern deep-learning methods, suffer from poor calibration [1] and [2], meaning that when the model is asked to predict with 80% confidence, we cannot expect it to be correct 80% of the time.
@@
Conformal prediction (CP) is a general framework for distribution-free uncertainty quantification. It can be used with essentially any machine learning algorithm and has guaranteed properties of validity, provided that the data are drawn from a probability distribution that is exchangeable, meaning essentially that any permutation is equally probable. For details on exchangeability, see [3]. The strongest notions of validity are achieved in the online setting, where Reality outputs successive examples $z_{n}:=(x_{n},y_{n})\in\boldsymbol{Z}:=\boldsymbol{X}\times\boldsymbol{Y}$, each of which consists of an object $x_{n}\in\boldsymbol{X}$ and its associated label $y_{n}\in\boldsymbol{Y}$. CP produces prediction sets, $\Gamma_{n}^{\varepsilon}$, at a user-specified significance level $\varepsilon$, using $z_{1},\dots,z_{n-1}$ and $x_{n}$ as inputs.
@@
If the exchangeability assumption is violated, the validity guarantees of CP are lost. Adaptive Conformal Inference (ACI) was suggested by Gibbs and Candès [4] as a method to produce prediction sets that are robust to violations of exchangeability. Fundamentally, ACI employs a simple feedback control mechanism that dynamically adjusts the working significance level based on recent prediction errors to steer the long-term empirical error rate towards a desired target $\varepsilon$. In this study, we show that there is no need to use ACI with a conformal predictor. The same guaranteed error rates hold even if we use the more general notion of a confidence predictor, which is often much more computationally tractable in the online setting. The applicability of ACI to non-conformal set predictors follows directly from its proof, which makes it somewhat surprising that, to the best of our knowledge, it has not been highlighted in the literature. Since nothing is gained by using the heavier machinery of CP in terms of error rate, the question is whether CP offers more efficient prediction sets.
We investigate this question empirically, both in the natural habitat of CP, the online setting, and the more popular inductive or batch setting, where inductive conformal predictors (ICP) are more computationally manageable. Our numerical experiments on publicly available datasets that are known to be non-exchangeable indicate that CP adds little beyond the computational effort, when compared with non-conformal confidence predictors (NCCP), when the assumption of exchangeability is violated in the online setting. In the more popular batch setting, we compare ICP with inductive NCCP (INCCP). One may suspect that sacrificing some of the available training data for calibration, which is required by ICP, may lead to less efficient (informally larger) prediction sets compared with alternative methods. This intuition is supported by our investigation of the efficiency over different calibration set sizes.
@@
The remainder of this paper is organised as follows. Section 2 discusses related work on the subject of uncertainty quantification in general, and CP for non-exchangeable data in particular. Section 3 introduces conformal and confidence predictors as well as the relevant notation. Readers familiar with the terminology used in [3] may find it convenient to skip ahead to Section 4, where ACI and its finite sample coverage guarantee is introduced, as well as a trivial but technically necessary extension of conformal and confidence predictors. Our main result is given in Section 5, where we restate Lemma 4.1 from [4], which is the key result used to prove the finite sample guarantee of ACI. We indicate that this lemma does not rely on any particular property of conformal, or even confidence predictors. We then argue in Section 5.1 that while not strictly necessary for the finite sample guarantee of ACI, confidence predictors represent the natural way of predicting at a confidence level. Numerical experiments on synthetic and real data are described in Section 6, and the results are presented in Section 7. Finally, Section 8 concludes the paper.
2 Related work
2.1 Uncertainty quantification
The main reference for CP is the book “Algorithmic Learning in a Random World” [5] and its second edition [3]. Several excellent shorter introductions to CP exist, including [6] and [7]. CP has been the subject of special issues of several journals, including Pattern Recognition [8] and Neurocomputing [9].
The most basic form of CP solves the uncertainty quantification problem by, instead of predicting a single point, outputting a prediction set associated with a user-specified significance level $\varepsilon$. Provided only that the data are exchangeable with no other distributional assumptions, the prediction sets will fail to contain the true label with probability $\varepsilon$, making conformal predictors provably well calibrated [3]. Depending on the specific use case, set predictions may not be desirable; in this case, the conformal framework can be used to produce probabilistic predictions both in classification (Venn predictors [10]) and regression (Conformal Predictive Systems [11]).
Bayesian methods [12] and [13] provide an alternative uncertainty quantification which, however, rely on a well-specified prior. Conformal predictors can be used in conjunction with Bayesian models for protection. For details, see [14] and the discussion of hard and soft models in [3].
It is well known [2] and [1] that many machine learning models are poorly calibrated, and several post-hoc calibration methods, including Platt scaling [15], temperature scaling [1] and Dirichlet scaling [16] exist to address this problem. Common for post-hoc calibration methods is to use a calibration set, not used in training, to learn a calibration map that transforms the model’s predictions to better calibrated ones.
2.2 Conformal prediction for non-exchangeable data
While powerful and general, the validity guarantees of CP rely on the exchangeability assumption, which is a limitation e.g. in time-series forecasting. To address these limitations, many authors have suggested solutions to retain at least some validity guarantees even when exchangeability is violated. These include introducing weights to attempt to deal with distribution shifts, such as [17] and [18], and what could be described as control theoretic approaches that attempt to achieve the desired error rate by varying the significance level.
ACI [4] falls into the second category of control theoretic approaches and has become a popular choice for non-exchangeable data. It has been suggested as a good method for time-series forecasting [19]. It was implemented in this mode in the Python package MAPIE [20]. In the time-series domain, it has also been suggested for multi-step-ahead forecasting by [21] and [22]. The more complicated conformal PID algorithm [23], inspired by classical proportional-integral-derivative (PID) control, can be viewed as a generalisation of ACI, as the authors recover it as a special case of conformal PID. Recent developments include adapting the step-size parameter in ACI [24] and [25].
@@ 3 Theoretical background
This section introduces the relevant theoretical background of mathematical objects and notation. Readers familiar with conformal prediction and the notation used in [3] may find it convenient to skip ahead to the next section, where ACI is introduced.
@@ 3.1 Confidence prediction in the online setting
called examples. For notational convenience, we write $z_{i}:=(x_{i},y_{i})\in\boldsymbol{X}\times\boldsymbol{Y}:=\boldsymbol{Z}$. The measurable space $\boldsymbol{Z}$ is called the example space. Thus, the infinite sequence (1) is an element of the measurable space $\boldsymbol{Z}^{\infty}$ and we assume that it is drawn from a probability distribution $P$ on $\boldsymbol{Z}^{\infty}$. The standard assumption in CP is that $P$ is exchangeable, but in this study, we make no such assumption, and $P$ can be any distribution.
Most machine learning methods are, so called, simple predictors, with the aim of predicting the label $y_{n}$. Here, we are interested in another type of prediction. Instead of predicting $y_{n}$, we want to predict subsets of $\boldsymbol{Y}$ of varying sizes, but large enough that we can be confident that $y_{n}$ will fall in them. Of course, this is a simple task; just predict $\boldsymbol{Y}$, and we are absolutely sure that the prediction set will contain $y_{n}$. However, we may be willing to accept a slightly smaller confidence level, provided that the prediction set is smaller and thus more informative.
@@
This last property is called the nested prediction set property.
Recall that $2^{\boldsymbol{Y}}$ is the set of all subsets of $\boldsymbol{Y}$. Thus, a confidence predictor outputs a subset of the label space based on all previous examples and the current object, and the idea is that it should contain the true label $y_{n}$ with a user-specified confidence. Moreover, the prediction sets are nested, as illustrated in Figure 1.
$\varepsilon=0.1$$\varepsilon=0.2$$\varepsilon=0.3$
Figure 1: Illustration of nested prediction sets.
Several machine-learning methods can be trivially turned into confidence predictors. In the classification case, confidence thresholding using confidence scores (for example, [26] or simple majority voting in ensemble methods are simple ways to define confidence predictors. For regression, many methods natively support confidence intervals. In the case of parametric methods, these are nested, such as ordinary least squares. Another method is quantile regression, but care must be taken, as many quantile regression methods do not guarantee that the prediction sets are nested. However, it is always possible to force prediction intervals to be nested by applying isotonic regression (see [3] or [27] for details) to the predicted quantiles and use the calibrated predictive quantiles to construct prediction intervals.
@@
at significance level $\varepsilon$. We say that $\Gamma$ makes an error in the $n$th trial if $y_{n}\notin\Gamma_{n}^{\varepsilon}$. More precisely,
@@ 3.2 Validity
is a sequence of independent Bernoulli random variables with parameter $\varepsilon$. In other words, the event of making an error is like getting heads when tossing a biased coin, where the probability of getting heads is always $\varepsilon$. There is also the notion of conservative validity, which is more complicated to state but essentially means that the error sequence is dominated by a sequence of independent Bernoulli variables with parameter $\varepsilon$. For the complete statement, we refer to [3].
@@ 3.3 Conformal predictors
Conformal predictors rely on the notion of a nonconformity score. Informally, this is a function that quantifies how “strange” or “unusual” an example $z$ is in relation to what we have seen before. More formally, given a sequence of examples $z_{1},\dots,z_{n}$, we form the bag (or multiset) $\Sigma_{n}:=\lbag z_{1},\dots,z_{n}\rbag$. The set of all bags of size $n$ formed from examples in $\boldsymbol{Z}$ is denoted $\boldsymbol{Z}^{(n)}$. A nonconformity measure is a measurable function
$A:\boldsymbol{Z}^{(*)}\times\boldsymbol{Z}$$\to\overline{\mathbb{R}}$ (10)
$(\Sigma,z)$ $\mapsto\alpha.$
Given a simple predictor, for example, a machine learning model that outputs the prediction $\hat{y}$ for the label $y$, a natural choice of the nonconformity measure is $\alpha:=|y-\hat{y}|$.
@@
The fraction |$\frac{|\{i=1,\dots,n:\alpha_{i}\geq\alpha_{n}\}|}{n}$ is called the p-value of the example $(x_{n},y)$.
@@
The main disadvantage of conformal predictors is their, often, intractable computational cost. In a regression setting where $\boldsymbol{Y}=\mathbb{R}$, it is theoretically necessary to compute the p-value in (11) for each real number, which is clearly impossible. For some special cases, such as ridge regression, efficient implementations exist, but generally the computational cost is too high. For this reason, inductive conformal predictors (ICP) have been introduced. We provide a brief description here and refer the interested reader to [3] for further details. Given a training set of size $l$, we split it into two parts: the proper training set $z_{1},\dots,z_{m}$ of size $m$ and the calibration set $z_{m+1},\dots,z_{l}$ of size $l-m$. For every test object $x_{i}$, compute the prediction set
@@
These are the most widely used conformal predictors in practice, but as always, there is a price to pay for the computational efficiency. ICPs are no longer exactly valid, but their validity property is of the probably approximately correct (PAC) type, with two parameters. For details on the training-conditional validity of ICPs, see [28].
Remark 1.
Vapnik [29] distinguished between induction and transduction, as applied to the prediction problem as follows. In inductive prediction, we start from some observed examples and extract some more or less general prediction rule, model or theory. This is the inductive step. For a test object, we derive a prediction from this general rule, which is the deductive step. Transductive prediction skips the inductive step and moves directly from observed examples to the prediction for a new object.
As definition 11 makes clear, conformal predictors are confidence predictors. From now on, we shall write non-conformal confidence predictors (NCCP) whenever we refer to confidence predictors that are not conformal predictors. An inductive confidence predictor that is not an inductive conformal predictor will be denoted by INCCP. The term confidence predictor refers to any confidence predictor, whether conformal or not.
@@ 4 Adaptive conformal inference
Adaptive conformal inference [4] has been suggested as a method to achieve asymptotic validity under non-exchangeable data. The idea is that instead of using the same significance level for all predictions, we use the online update
@@
where $\gamma$ is the step size (or learning rate). In other words, if we made an error in the last step, we increase the significance level; otherwise, we decrease it. Gibbs and Candès proved the following finite sample guarantees of ACI.
@@
which in particular converges to 0 as $N\to\infty$, ensuring asymptotic validity. Note that if we want a certain absolute error deviation bound, $\delta>\frac{\max\{\varepsilon_{1},1-\varepsilon_{1}\}+1}{N}$, for a desired error rate $\varepsilon$, with a finite sample size $N$, we can almost surely achieve it by choosing
@@
The update rule (15) resembles a simple controller that increases or decreases the significance level to counteract deviations from the target error rate, $\varepsilon$. Importantly, the iteration (15) can cause $\varepsilon_{n}\leq 0$ or $\varepsilon_{n}\geq 1$. Technically, prediction sets are undefined for confidence predictors for $\varepsilon\notin(0,1)$. Thus, we introduce a trivial extension that is compatible with ACI.
@@
Since the difference between confidence predictors and extended confidence predictors is minor, and since any confidence predictor can be trivially modified to be an extended confidence predictor, we will use the terms interchangeably.
5 The use of conformal predictors is not essential
Interestingly, the proof of (16) does not use any property of a conformal, or even confidence predictor. In fact, all that is required of a predictor $\Gamma$ is that $\emptyset$ is predicted for $\varepsilon\geq 1$ and $\boldsymbol{Y}$ for $\varepsilon\leq 0$. For all other significance levels, even completely random prediction sets ensure (16). The result is proved by Lemma 4.1 in [4], which we reproduce here with a notation slightly modified to align with ours.